OpenStudy (anonymous):

Check if its right: A school auditorium can hold 386 people. For a choir concert there will be 23 staff in attendance. Which inequality describes the number of people each of the 33 choir members can invite to fit the auditorium. THe choices and my answer will be posted below in one minute.

OpenStudy (anonymous):

\[choice 1: g \ge11 guests\] \[choice 2: g \le 11 guests\] \[choice 3: g \ge 12 guests\] \[choice 4: g \le 12 guests\] I chose choice b is it right?

OpenStudy (anonymous):

yes it's right @stylestar5678

OpenStudy (anonymous):

do u know how to get the answer? @stylestar5678

OpenStudy (anonymous):

yes i do i just wasnt really sure if i got the symbol right. but i figured it out thanks